OpenStudy (anonymous):

A parent increases a child's allowance by 15% each year. If the allowance is $3 now, when will it reach $15?

OpenStudy (anonymous):

t = no. of years for this to happen : 3*(1.15^t) = 15 divide both sides by 3 1.15^t = 1.15^t = 5 : use logs here log(1.15^t) = log(5) which is t*log(1.15) = log(5) : t =log5/log1.15 ; enter into your calc t ~ 11.5 years

OpenStudy (paxpolaris):

15% increrase means: in one year you will have \((100\%+15\%)=115\%\) of the original amount 115%=1.15